OpenStudy (anonymous):

does anyone know calc AB? evaluate: lim as x approaches 2 4/radical 18-x -1 divided by x-2/radical 18-x

OpenStudy (anonymous):

@loser66 please help!!!! youre my only option test tomorrow

OpenStudy (loser66):

can you use the draw box to draw it out? or use latex ?? It is unclear to me

OpenStudy (anonymous):

https://docs.google.com/viewer?a=v&pid=sites&srcid=ZGVmYXVsdGRvbWFpbnxtYmFybmVzbWF0aHxneDozMjBiMGIyMGEwZGMwYzc @Loser66 number 22 in case I wrote it weird

OpenStudy (anonymous):

know it, yes, read it, no

OpenStudy (anonymous):

@satellite73 @Loser66 I posted a link to the review question im asking

OpenStudy (loser66):

Let satelite73 help. :)

OpenStudy (anonymous):

bunch of algebra, eliminate the compound fraction you might want to start by multplying top and bottom by \(\sqrt{18-x}\)

OpenStudy (anonymous):

done

OpenStudy (anonymous):

\[\frac{4-\sqrt{18-x}}{x-2}\] maybe?

OpenStudy (loser66):

yup, then l'hopital rule

OpenStudy (anonymous):

I think that's it

OpenStudy (anonymous):

then rationalize the numerator by multiplying top and bottom by \(4+\sqrt{18-x}\)

OpenStudy (anonymous):

yes I remember now!!!! thanks sooo sooo much

OpenStudy (anonymous):

no need for l'hopital and besides you didn't get there yet

OpenStudy (anonymous):

ok. so I got 16- radical18-x over (x-2)(4+rad18-x)

OpenStudy (loser66):

why not?@satellite73 still get the form 0/0

OpenStudy (anonymous):

sorry the radicals cancel out on the top foregot

OpenStudy (anonymous):

@loser can you help me he left?

OpenStudy (anonymous):

\[(-x-2)\div(x-2)(4+\sqrt{18-x}\]

OpenStudy (anonymous):

@loser66 shouldn't (-x-2) be (x-2)

OpenStudy (anonymous):

that's wrong. The correct answer is 1/8 I have all fo the answers just no work. The way I did it was correct I just got -x-2 instead of x-2 so those two would cancel out. Once those two cancel out you plug two into the equation and get 1/8 :)

OpenStudy (anonymous):

Im in ab all this chapter is about ar elimits derivitives are next chapter.

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

K

OpenStudy (loser66):

ok, so take his way, :)

OpenStudy (loser66):

\[lim_{x\rightarrow 2}\dfrac{(4-\sqrt{18-x})(4+\sqrt{18-x})}{(x-2)(4+\sqrt{18-x})}\]

OpenStudy (anonymous):

yup did that but im getting -x-2 instead of x-2 do you see where I am going wrong?

OpenStudy (loser66):

numerator: = 4^2 -(18-x) = 16-18+x= x-2

OpenStudy (anonymous):

thanks! can you help me with another?

OpenStudy (loser66):

cancel out with the denominator, you just have \[lim_{x\rightarrow 2}\dfrac{1}{(4+\sqrt{18-x})}=1/8\]

OpenStudy (loser66):

you do, I check

OpenStudy (anonymous):

ok. ill start a new topic. im gong to do number 10